# A Chinese Continuous Sign Language Dataset Based on Complex Environments

### v1

@@ III-A CE-CSL dataset
1 A {CJK*}UTF8gbsn今晚瑜伽课什么时候开始呢? {CJK*}UTF8gbsn今天/晚上/瑜伽/课/开始/什么/时间/? {CJK*}UTF8gbsn瑜伽
2 A {CJK*}UTF8gbsn一大盆异常茂盛的法国吊兰。 {CJK*}UTF8gbsn大/茂盛/法国/花/名字/吊兰/。 {CJK*}UTF8gbsn吊兰
3 A {CJK*}UTF8gbsn人生有许多困难和失败。 {CJK*}UTF8gbsn人生/困难/失败/多/。 {CJK*}UTF8gbsn人生
4 A {CJK*}UTF8gbsn他送给我们一件他旅游的纪念品。 {CJK*}UTF8gbsn他/旅游/纪念//送 (我)/。 {CJK*}UTF8gbsn
@@
Transcription Standards for Converting Chinese Sentences into Sign Language Sentences Detailed Explanation
When a single word in the dictionary has multiple sign language representations, the specific representation used should be noted at the end of the word. For example: For the word ”{CJK*}UTF8gbsn事情” (thing), if there are two sign language representations in the dictionary—”{CJK*}UTF8gbsn事情1” (thing1) and ”{CJK*}UTF8gbsn事情2” (thing2)—it is necessary to specify which representation is used during transcription.
Select the corresponding words based on the actual meaning conveyed by the sentence. For example: The term ”{CJK*}UTF8gbsn不会” has two meanings: ”do not know” and ”impossible.” However, if the dictionary only provides a sign for ”do not know,” during the transcription process, you should choose the signs for ”impossible” or ”cannot” based on the actual meaning conveyed by the sentence.
For terms that involve only the change of numbers, it is assumed that the dictionary contains the term, and no special annotation is required. For example: From January to December, and from Monday to Sunday, the dictionary may only include entries for January and Monday, but if it states ”and so on,” it is assumed that the dictionary includes representations for all months and days of the week. Therefore, no special annotation is needed for the signs of other months or days.
For words that do not exist in the dictionary, use regional sign language representations. For example: If the word ”{CJK*}UTF8gbsn英语” (English) does not exist in the dictionary, then use the regional sign language representation. Words that use regional sign language representations will be specified in the notes.
Directional words should be followed by ( ) to specify the subject of expression. For example: ”{CJK*}UTF8gbsn他帮我开门” is transcribed as ”{CJK*}UTF8gbsn他/帮(我)/开门.”
For body turns, add [ ] after the action description words to specify the direction. For example: ”{CJK*}UTF8gbsn我想在饮料里放些冰块” is transcribed as ”{CJK*}UTF8gbsn喝[左]/冰[右]/块/抓[右]/放[左].”
@@ III-B Dataset annotation
As a visual language, sign language carries unique grammatical structures and lexical ordering that differ significantly from traditional written or spoken languages, making direct mapping challenging. Therefore, the CE-CSL dataset introduces deep involvement from professional sign language translators, who are responsible for converting sign glosses into video content. Additionally, the dataset provides spoken language translations as a reference to enhance multidimensional understanding. Notably, we have conducted meticulous word segmentation for the sign glosses, whereas the spoken language translations maintain complete sentence structures without segmentation.
@@
The construction of the CE-CSL dataset relies on 12 professional sign language performers, including both professional sign language interpreters and hearing-impaired individuals, who jointly undertake the task of translating spoken language into sign glosses. This process involves a deep understanding of spoken language sentences, followed by their reformation into expressions that conform to sign language grammar, which are then solidified through video demonstrations. All videos and their corresponding annotations are subject to comprehensive review and correction by experienced sign language interpreters after individual tasks are completed, ensuring the accuracy and consistency of the dataset.
@@
The current sign language environment in China exists in a state where both standard sign language and regional sign languages coexist. Although efforts have been made to promote a standard sign language dictionary (referred to as the dictionary), a unified standard has yet to be fully established. Against this backdrop, the CE-CSL dataset adopts an inclusive approach. During the transcription process, in addition to relying on standard sign language, regional sign languages are appropriately integrated, especially when there is no corresponding sign in the standard dictionary for a spoken word. For words that use regional sign languages, we explicitly label them in the annotations, as shown in Table IV, to ensure the transparency and traceability of the dataset.
@@
The transcription of sign language is not a mechanical process but one guided by a series of norms. For instance, when encountering words that exist in written or spoken language but are missing from the sign language dictionary, we follow the intent of the sentence and select the most appropriate substitute words. When faced with multiple sign language expressions for synonyms listed in the dictionary, we append specific gesture versions after the words to ensure accuracy, as listed in Table V. These guidelines ensure the rigor and practicality of the CE-CSL dataset, providing researchers with reliable material for their studies.
@@
In summary, the CE-CSL dataset not only demonstrates the richness of sign language videos in quantity but also exhibits professionalism and meticulous annotation and transcription processes, providing a solid data foundation for the advancement of sign language recognition technology. By integrating both standard and regional sign languages, along with rigorous transcription standards, it effectively propels the field of sign language recognition forward.
@@ III-C Methodology
This paper proposes a TFNet for CSLR, as illustrated in Figure 3. TFNet primarily consists of two key components: a frame-level feature extraction module and a sequence feature extraction module. Initially, in the frame-level feature extraction stage, we employ a CNN to efficiently extract feature information from frames of sign language videos. Subsequently, in the sequence feature extraction part, the model is designed with two independent yet complementary branches. In one branch, we model the temporal sequence using a combination of 1D CNN and Bidirectional Long Short-Term Memory Networks (Bi-LSTM)[31] to capture the temporal dynamics of sign language actions. In the other branch, we transform the temporal features into the frequency domain via Discrete Fourier Transform (DFT)[32] and similarly model the frequency-domain sequence using 1D CNN and Bi-LSTM to obtain sequence features within the frequency domain. Ultimately, the outputs from these two branches are merged and subjected to a classifier for joint prediction, thereby achieving effective recognition of sign language sentences.
1) TFNet: CSLR involves translating continuous sign language videos into written phrases or spoken language understandable by the general population. Let the size of the input continuous sign language video frames be $T$, then the continuous sign language video $V=(x_{1},x_{2},...,x_{T})=\left\{x_{t}\left|{}_{1}^{T}\in\mathbb{R}^{T\times C\times H\times W}\right.\right\}$, where $x_{t}$ is the image of the t-th frame in the video, $H\times W$ is the dimension of $x_{t}$, and C is the number of channels. In the TFNet model, we first extract frame-level features and then extract time-frequency sequence features. Let the frame-level feature extractor be $F_{f}$, yielding the feature representation as follows:
@@
where $f_{frame}$ is the frame-level feature after extraction, and $C^{\prime}$ is the number of channels after frame-level feature extraction. Then, the frame-level feature $f_{frame}$ is input into two branches. In one branch, $f_{frame}$ is directly fed into the temporal sequence feature extractor $F_{\mathrm{sequences}_{t}}$. In the other branch, $f_{frame}$ is first transformed via discrete Fourier transform to convert the temporal data into frequency domain data, which is then input into the frequency sequence feature extractor $F_{\mathrm{sequences}_{f}}$. The methods used by these two sequence feature extractors are the same (both being 1D CNN + Bi-LSTM), leading to the following:
$f_{temporal}=F_{\mathrm{sequences}_{t}}(f_{frame})\in\mathbb{R}^{T^{\prime}\times C^{\prime\prime}}$ (2)
$f_{freq}=F_{\mathrm{sequences}_{f}}(DFT(f_{frame}))\in\mathbb{R}^{T^{\prime}\times C^{\prime\prime}}$ (3)
Among them, $f_{temporal}$ represents the temporal domain sequence features after extraction, and $f_{freq}$ represents the spectral domain sequence features after extraction. $T^{\prime}$ is the length of the sequence features after downsampling in the time dimension, and $C^{\prime\prime}$ is the number of channels after temporal sequence feature extraction. The size of the spectral domain sequence features is consistent with that of the temporal domain sequence features. Finally, the temporal domain sequence features $f_{temporal}$ and the spectral domain sequence features $f_{freq}$ are summed and passed through a fully connected layer for recognition:
@@
Among them, $f_{classification}$ represents the final recognition result, and $l$ is the number of words.
2) Loss Function: The overall loss function $L_{sum}$ of the TFNet model consists of three parts: two auxiliary VAE loss functions[33], $L_{VAE_{t}}$ and $L_{VAE_{f}}$, and a final CTC (Connectionist Temporal Classification) loss function[34]. Specifically, $L_{VAE_{t}}$ denotes the VAE loss function applied to the temporal domain sequence features, and $L_{VAE_{f}}$ denotes the VAE loss function applied to the spectral domain sequence features. The final CTC loss function is used to calculate the difference between the predicted sequence and the true labels, guiding the model to learn the correct sign language recognition path. Therefore, $L_{sum}$ is defined as follows:
@@
TABLE VI: Performance comparison with other advanced methods on RWTH, RWTH-T, and CSL-Daily.
Methods RWTH RWTH-T CSL-Daily
Dev(%) Test(%) Dev(%) Test(%) Dev(%) Test(%)
VAC[33] 21.2 22.3 - - - -
MSTNet[30] 20.3 21.4 - - - -
SEN[35] 19.5 21.0 19.3 20.7 31.1 30.7
AdaSize[36] 19.7 20.9 19.7 21.2 31.3 30.9
STMC[37] 21.1 20.7 19.6 21.0 - -
STENet[38] 19.3 20.3 19.4 21.1 28.9 28.9
HST-GNN[39] 19.5 19.8 20.1 20.3 - -
DSE[40] 18.6 19.8 18.9 19.9 28.6 27.6
TB-Net[41] 18.9 19.6 18.8 20.0 28.4 28.2
CorrNet[26] 18.8 19.4 18.9 20.5 30.6 30.1
TCatet[42] 18.1 18.9 18.3 19.4 29.7 29.3
MAM-FSD[28] 19.2 18.8 18.2 19.4 25.8 24.5
TwoStream-SLR[43] 18.4 18.8 17.7 19.3 25.4 25.3
TENet 18.7 18.6 18.0 19.1 25.1 23.5
TABLE VII: We conduct experiments on the CE-CSL dataset using the state-of-the-art continuous sign language recognition model using WER as an evaluation criterion, with lower WER indicating better performance.
Methods CE-CSL
Dev(%) Test(%)
MSTNet[30] 54.4 53.0
CorrNet[26] 47.2 46.5
SEN[35] 46.5 45.3
VAC[33] 45.1 43.3
MAM-FSD[28] 44.9 44.7
THNet 42.1 41.9
IV Experiment
IV-A Dataset and judgment criteria
In addition to the CE-CSL dataset constructed in this paper, we conducted experiments on three other large-scale publicly available datasets, including RWTH[11], RWTH-T[12], and CSL-Daily[16], aimed at verifying the wide applicability and effectiveness of the proposed method. The RWTH dataset contains a total of 6841 different video clips, which is derived from sign language videos recorded by the German Weather Radio and Television (DWRT). The official pre-defined splits consist of a training set (5,672 videos), a validation set (540 videos), and a test set (629 videos). The RWTH-T dataset is an extended version of the RWTH dataset with an expanded size, where the dataset is subdivided into 7,096 training videos, 519 validation videos, and 642 test videos. The CSL-Daily dataset is a large Chinese sign language dataset with an annotated vocabulary of 2,000 words and a Chinese text vocabulary of 2,343 words. The dataset consists of 18,401 training samples, 1,077 validation samples, and 1,176 test samples.
$WER=100\%\times\frac{ins+del+sub}{sum}$ (6)
where ins represents the number of words to be inserted, $del$ represents the number of words to be deleted, $sub$ represents the number of words to be replaced, and $sum$ represents the total number of words in the label.
IV-B Implementation rules
TABLE VIII: Qualitative results for the CE-CSL dataset, with red font indicating identification errors.
@@ IV-B Implementation rules
Original sentence {CJK*}UTF8gbsn他从小就相当律师。 - {CJK*}UTF8gbsn我可以支持你去运动。 -
Gloss ground truth {CJK*}UTF8gbsn他/小孩时间/开始/做/律师/希望/。 - {CJK*}UTF8gbsn我/可以支持/你/去/运动/。 -
Pred. (SEN) {CJK*}UTF8gbsn他/时间/开始/。 50.0 {CJK*}UTF8gbsn我/可以你/经济/木头/。 42.9
Pred. (CorrNet) {CJK*}UTF8gbsn他时间/开始/做/希望/。 25.0 {CJK*}UTF8gbsn我/可以/你/。 42.9
Pred. (VAC) {CJK*}UTF8gbsn他/小孩时间/开始/工作/困难/。 37.5 {CJK*}UTF8gbsn我/可以/你//锻炼/。 42.9
Pred. (MAM-FSD) {CJK*}UTF8gbsn他/小孩时间/开始/做/律师/希望/。 0.0 {CJK*}UTF8gbsn我/可以/支持/你/锻炼/。 28.6
Pred (TFNet) {CJK*}UTF8gbsn他/小孩时间/开始/做/律师/希望/。 0.0 {CJK*}UTF8gbsn我/可以/支持/你/去/运动/。 0.0
IV-C Experimental results
The TFNet model is tested on RWTH, RWTH-T, CSL-Daily, and CE-CSL datasets, with the model recognition accuracies presented in Tables VI and VII.
TABLE IX: Ablation study of different domain sequence features.
@@ IV-C Experimental results
TABLE X: Ablation study of different sequence feature extractors (original sequence feature extractors with different models retained).
@@
TABLE XI: Ablation study of different loss combinations.

### v2

@@ III-A CE-CNSL dataset
1 A 今晚瑜伽课什么时候开始呢? 今天/晚上/瑜伽/课/开始/什么/时间/? 瑜伽
2 A 一大盆异常茂盛的法国吊兰。 大/茂盛/法国/花/名字/吊兰/。 吊兰
3 A 人生有许多困难和失败。 人生/困难/失败/多/。 人生
4 A 他送给我们一件他旅游的纪念品。 他/旅游/纪念//送 (我)/。
The CE-CNSL dataset encompasses 12 sign language performers (labeled A through L), including eight females and four males. Notably, performers I and J are hearing-impaired individuals, while the remaining performers are professional sign language interpreters, ensuring a diverse and authentic representation of sign language expressions within the dataset. The dataset comprises a total of 5,988 continuous sign language videos. According to strict division ratios, there are 4,973 sign language videos in the training set, 515 in the validation set, and 500 in the test set. These videos collectively involve 3,515 Chinese words, covering a broad range of daily communication needs. The video lengths vary widely, ranging from the shortest at 39 frames to the longest at 530 frames, showcasing the richness and flexibility of sign language expression. Detailed statistics of the CE-CNSL dataset are presented in Table II.
We also provide a comprehensive listing of key parameters for each performer in Table III, including the recording equipment used, the number of videos, age, and gender, offering detailed baseline information for future research. Given the open setting for background environments, the CE-CNSL dataset includes over 70 different living scenarios, greatly enriching the diversity of the data. To better understand the diversity of backgrounds in the CE-CNSL dataset, we present a series of representative scenes in Figure 2, including indoor, outdoor, park, street, mall, office, and others, all sourced from our dataset. Through these images, one can clearly see the extensive range of scenarios covered by the CE-CNSL dataset, further demonstrating its valuable contribution as a research resource and its significance in advancing the practical application of CSL recognition technology.
@@ 4 A 他送给我们一件他旅游的纪念品。 他/旅游/纪念//送 (我)/。
Transcription Standards for Converting Chinese Sentences
into Sign Language Sentences
Detailed Explanation
When a single word in the dictionary has multiple sign
language representations, the specific representation
used should be noted at the end of the word.
For example: For the word ”事情” (thing), if there are
two sign language representations in the dictionary—”事情1”
(thing1) and ”事情2” (thing2)—it is necessary to specify which
representation is used during transcription.
Select the corresponding words based on the actual meaning
conveyed by the sentence.
For example: The term ”不会” has two meanings: ”do not know”
and ”impossible.” However, if the dictionary only provides a sign
for ”do not know,” during the transcription process, you should
choose the signs for ”不可能(impossible)” or ”不行(cannot)” based on the
actual meaning conveyed by the sentence.
For terms that involve only the change of numbers, it is
assumed that the dictionary contains the term, and no special
annotation is required.
For example: From January to December, and from Monday to
Sunday, the dictionary may only include entries for January and
Monday, but if it states ”and so on,” it is assumed that the
dictionary includes representations for all months and days of the
week. Therefore, no special annotation is needed for the signs of
other months or days.
For words that do not exist in the dictionary, use regional
sign language representations.
For example: If the word ”英语” (English) does not exist
in the dictionary, then use the regional sign language
representation. Words that use regional sign language
representations will be specified in the notes.
Directional words should be followed by ( ) to specify the
subject of expression.
For example: ”他帮我开门” is transcribed as ”他/帮(我)/开门.”
For body turns, add [ ] after the action description words
to specify the direction.
For example: ”我想在饮料里放些冰块” is transcribed
as ”喝[左]/冰[右]/块/抓[右]/放[左].”
@@ III-B Dataset annotation
As a visual language, sign language carries unique grammatical structures and lexical ordering that differ significantly from traditional written or spoken languages, making direct mapping challenging. Therefore, the CE-CNSL dataset introduces deep involvement from professional sign language translators, who are responsible for converting sign glosses into video content. Additionally, the dataset provides spoken language translations as a reference to enhance multidimensional understanding. Notably, we have conducted meticulous word segmentation for the sign glosses, whereas the spoken language translations maintain complete sentence structures without segmentation.
@@
The CE-CNSL dataset revolves around topics such as family life, learning, shopping, and daily communication. The sentences are derived from daily communication statements, weather forecasts, news content, and online articles. Regarding the collection process of the dataset, firstly, we collected a large number of written or colloquial sentences from different sources as raw data. Then, we deduplicated the raw data, leaving unique sentences as the final target sentences. Then, 12 professional sign language presenters, including professional sign language translators and hearing-impaired individuals, jointly undertook the task of converting from oral translation to symbol gloss. This process involves a deep understanding of spoken language sentences, followed by their reformation into expressions that conform to sign language grammar, which are then solidified through video demonstrations. All videos and their corresponding annotations are subject to comprehensive review and correction by experienced sign language interpreters after individual tasks are completed, ensuring the accuracy and consistency of the dataset.
@@
The current sign language environment in China exists in a state where both standard sign language and regional sign languages coexist. Although efforts have been made to promote a standard sign language dictionary (referred to as the dictionary), a unified standard has yet to be fully established. Against this backdrop, the CE-CNSL dataset adopts an inclusive approach. During the transcription process, in addition to relying on standard sign language, regional sign languages are appropriately integrated, especially when there is no corresponding sign in the standard dictionary for a spoken word. For words that use regional sign languages, we explicitly label them in the annotations, as shown in Table IV, to ensure the transparency and traceability of the dataset.
@@
The transcription of sign language is not a mechanical process but one guided by a series of norms. For instance, when encountering words that exist in written or spoken language but are missing from the sign language dictionary, we follow the intent of the sentence and select the most appropriate substitute words. When faced with multiple sign language expressions for synonyms listed in the dictionary, we append specific gesture versions after the words to ensure accuracy, as listed in Table V. These guidelines ensure the rigor and practicality of the CE-CNSL dataset, providing researchers with reliable material for their studies.
@@
In summary, the CE-CNSL dataset not only demonstrates the richness of sign language videos in quantity but also exhibits professionalism and meticulous annotation and transcription processes, providing a solid data foundation for the advancement of sign language recognition technology. By integrating both standard and regional sign languages, along with rigorous transcription standards, it effectively propels the field of sign language recognition forward.
@@ III-C Methodology
This paper proposes a TFNet for CSLR, as illustrated in Figure 3. TFNet primarily consists of two key components: a frame-level feature extraction module and a sequence feature extraction module. Initially, in the frame-level feature extraction stage, we employ a CNN to efficiently extract feature information from frames of sign language videos. Subsequently, in the sequence feature extraction part, the model is designed with two independent yet complementary branches. In one branch, we model the temporal sequence using a combination of 1DCNN and Bidirectional Long Short Term Memory Networks (BiLSTM)[30] to capture the temporal dynamics of sign language actions. In the other branch, we transform the temporal features into the frequency domain via Discrete Fourier Transform (DFT)[31] and similarly model the frequency-domain sequence using 1DCNN and BiLSTM to obtain sequence features within the frequency domain. Ultimately, the outputs from these two branches are merged and subjected to a classifier for joint prediction, thereby achieving effective recognition of sign language sentences.
[…]
@@
where $f_{frame}$ is the frame-level feature after extraction, and $C^{\prime}$ is the number of channels after frame-level feature extraction. Then, the frame level feature $f_{frame}$ is input into the sequence feature extractor, which includes two branch structures: one branch focuses on parsing the dynamic characteristics of sign language actions over time based on time-domain feature sequences; Another branch analyzes the dynamic characteristics of sign language actions over time from a frequency domain perspective based on frequency domain feature sequences. This article uses the discrete Fourier transform to transform the eigenvectors into the frequency domain. The following is the basic formula of the discrete Fourier transform:
$Z[k]=\sum_{n=0}^{N-1}z[n]\cdotp e^{-j2\pi kn/N}$ (2)
Among them, $z[n]$ is the input discrete time series, $Z[k]$ is the frequency domain representation of the output, $k=0,1,2,\ldots,N-1$. $j$ is the imaginary unit that satisfies $j^{2}=-1$. $e^{-j2\pi kn/N}$ is a complex exponential function used to calculate the coefficients of each frequency component. Substitute $f_{frame}$ into formula (2) to calculate the frequency domain sequence feature, and let the discrete Fourier transform be $DFT(\cdot)$, then,
$f^{{}^{\prime}}_{frame}=DFT(f_{frame})\in\mathbb{R}^{T\times C^{\prime}}$ (3)
TABLE VI: Performance comparison with other advanced methods on RWTH, RWTH-T, CSL-Daily, and CE-CNSL.
Methods RWTH RWTH-T CSL-Daily CE-CNSL
Dev(%) Test(%) Dev(%) Test(%) Dev(%) Test(%) Dev(%) Test(%)
VAC[32] 21.2 22.3 - - - - 45.1 43.3
MSTNet[24] 20.3 21.4 - - - - 54.4 53.0
SEN[33] 19.5 21.0 19.3 20.7 31.1 30.7 46.5 45.3
AdaSize[34] 19.7 20.9 19.7 21.2 31.3 30.9 - -
STMC[25] 21.1 20.7 19.6 21.0 - -
STENet[35] 19.3 20.3 19.4 21.1 28.9 28.9 - -
HST-GNN[36] 19.5 19.8 20.1 20.3 - - - -
DSE[37] 18.6 19.8 18.9 19.9 28.6 27.6 - -
TB-Net[38] 18.9 19.6 18.8 20.0 28.4 28.2 - -
CorrNet[22] 18.8 19.4 18.9 20.5 30.6 30.1 47.2 46.5
TCNet[39] 18.1 18.9 18.3 19.4 29.7 29.3 - -
MAM-FSD[23] 19.2 18.8 18.2 19.4 25.8 24.5 44.9 44.7
TwoStream-SLR[40] 18.4 18.8 17.7 19.3 25.4 25.3 - -
TFNet 18.7 18.6 18.0 19.1 25.1 23.5 42.1 41.9
where $f^{{}^{\prime}}_{frame}$ is the result of the discrete Fourier transform of the temporal feature sequence.
One branch directly inputs $f_{frame}$ into the time-domain sequence feature extractor, which is composed of a hybrid neural network consisting of 1DCNN and BiLSTM. Firstly, $f_{frame}$ is transposed and passed through 1DCNN, where each convolution is followed by a BN layer, ReLU activation function, and max pooling layer.
$\mathrm{g}_{i}=\mathrm{MaxPool}(\mathrm{ReLU}(\mathrm{BN}(\sum_{k=1}^{K}\left(f_{i+k-1}\cdot w_{k}\right)+b)))$ (4)
Among them, $w$ is the weight of the convolution kernel, $K$ is the number of convolution kernels, $b$ is the bias term, $g_{i}$ is the i-th output data, $f_{i}$ is the i-th data of the $f_{frame}$ feature vector, and $i\in[0,T]$. All output data are defined as $f_{gloss}\in\mathbb{R}^{C^{\prime\prime}\times T^{\prime}}$, where $T^{\prime}$ is the length of the downsampled sequence features in the time dimension, and $C^{\prime\prime}$ is the size of the number of channels extracted from the temporal features. Then transpose it and output it to BiLSTM.
$f_{temporal}=BiLSTM\left(f_{gloss}\right)\in\mathbb{R}^{T^{\prime}\times C^{\prime\prime}}$ (5)
Among them, $f_{temporal}$ is the time-domain sequence feature extracted from the sequence feature. Another branch inputs $f^{{}^{\prime}}_{frame}$ into the frequency domain sequence feature extractor $F_{\mathrm{sequences}_{f}}$, and the methods used by these two sequence feature extractors are the same (both 1DCNN+BiLSTM), which yields,
$f_{freq}=F_{\mathrm{sequences}_{f}}(f^{{}^{\prime}}_{frame})\in\mathbb{R}^{T^{\prime}\times C^{\prime\prime}}$ (6)
Among them, $f_{freq}$ is the extracted frequency domain sequence feature, and the size of the frequency domain sequence feature is consistent with that of the time domain sequence feature. Finally, the time-domain sequence feature $f_{temporal}$ and frequency-domain sequence feature $f_{freq}$ are summed and identified through a fully connected layer.
@@
Where $f_{classification}$ is the final recognition result, and $l$ is the number of words.
2) Loss Function: The overall loss function $L_{sum}$ of the TFNet model consists of three parts: two auxiliary VAE loss functions , $L_{VAE_{t}}$ and $L_{VAE_{f}}$, and a final CTC (connectionist temporary classification)[41] loss function. The VAE loss function was proposed by Min et al.[32] in 2021, which includes two parts: visual enhancement and visual alignment loss.
(1) Visual Enhancement Loss
In order to enhance the feature extraction capability of the feature extractor, researchers have implemented it by providing an auxiliary loss to the feature extractor, which is the Visual Enhancement (VE) loss. This auxiliary loss is designed to allow the feature extractor to make predictions based on local visual information, which significantly improves the model’s ability to capture details compared to traditional methods. Specifically, a CTC loss is introduced as a VE loss on the auxiliary classifier in the feature extraction stage with the formula:
$L_{VE}=L_{CTC}=-\log p(l\mid X;\theta)$ (8)
where $l$ denotes the label, $X$ denotes the input sequence, and $\theta$ denotes the model parameters.
(2) Visual alignment loss
Since the VE loss function lacks contextual information and is independent of the main loss, this can lead to misalignment between the two classifiers, which affects the overall performance of the model. To circumvent this problem, the researchers used a Visual Alignment (VA) loss function for optimization.The design of the VA loss function is based on the idea of knowledge distillation, which treats the whole network as a teacher model and the visual feature extractors as student models. In this way, the student model can learn more contextual information and deeper feature representations from the teacher model, thus improving the accuracy and robustness of its prediction. And the high temperature parameter $\tau$ is introduced to soften the probability distribution of the peak response when calculating the VA loss. This softening operation can make the probability distribution smoother, thus reducing the risk of overfitting.
$L_{VA}=KL\left(\mathrm{softmax}(\frac{Z}{\tau}),\mathrm{softmax}(\frac{\tilde{Z}}{\tau})\right)$ (9)
where $Z$ denotes the predicted value and $\tilde{Z}$ denotes the standard output.
Specifically, $L_{VAE_{t}}$ is dedicated to applying the VAE loss function to the time-domain sequence features, which aims to achieve a sharp capture and efficient representation of the temporal dynamics; meanwhile, $L_{VAE_{f}}$ focuses on executing the VAE loss function to the frequency-domain sequence features, which aims to strengthen the model’s effectiveness in processing the frequency-domain information. The final CTC loss function, which is used to calculate the difference between the predicted sequences and the real labels, guides the model to learn the correct sign language recognition path, then $L_{sum}$ is defined as:
@@
TABLE VII: Qualitative results for the CE-CNSL dataset, with red font indicating identification errors.
@@
Original sentence 他从小就相当律师。 - 我可以支持你去运动。 -
Gloss ground truth 他/小孩时间/开始/做/律师/希望/。 - 我/可以支持/你/去/运动/。 -
Pred. (SEN) 他/时间/开始/。 50.0 我/可以你/经济/木头/。 42.9
Pred. (CorrNet) 他时间/开始/做/希望/。 25.0 我/可以/你/。 42.9
Pred. (VAC) 他/小孩时间/开始/工作/困难/。 37.5 我/可以/你//锻炼/。 42.9
Pred. (MAM-FSD) 他/小孩时间/开始/做/律师/希望/。 0.0 我/可以/支持/你/锻炼/。 28.6
Pred (TFNet) 他/小孩时间/开始/做/律师/希望/。 0.0 我/可以/支持/你/去/运动/。 0.0
TABLE VIII: Ablation study of different domain sequence features.
@@
TABLE IX: Ablation study of different sequence feature extractors (original sequence feature extractors with different models retrained).
@@
TABLE X: Ablation study of different loss combinations.
@@
IV Experiment
IV-A Dataset and judgment criteria
In addition to the CE-CNSL dataset constructed in this paper, we conducted experiments on three other large-scale publicly available datasets, including RWTH[8], RWTH-T[9], and CSL-Daily[11], aimed at verifying the wide applicability and effectiveness of the proposed method. The RWTH dataset contains a total of 6841 different video clips, which is derived from sign language videos recorded by the German Weather Radio and Television (DWRT). The official pre-defined splits consist of a training set (5,672 videos), a validation set (540 videos), and a test set (629 videos). The RWTH-T dataset is an extended version of the RWTH dataset with an expanded size, where the dataset is subdivided into 7,096 training videos, 519 validation videos, and 642 test videos. The CSL-Daily dataset is a large Chinese sign language dataset with an annotated vocabulary of 2,000 words and a Chinese text vocabulary of 2,343 words. The dataset consists of 18,401 training samples, 1,077 validation samples, and 1,176 test samples.
$WER=100\%\times\frac{ins+del+sub}{sum}$ (11)
where insins represents the number of words to be inserted, deldel represents the number of words to be deleted, subsub represents the number of words to be replaced, and sumsum represents the total number of words in the label.
IV-B Implementation rules
IV-C Experimental results
The TFNet model is tested on RWTH, RWTH-T, CSL-Daily, and CE-CNSL datasets, with the model recognition accuracies presented in Tables VI and VII.